Celcius temperature readings can be converted to Fahrenheit readings using the formula F=9/5C+32. What is the Fahrenheit temperature that cooresponds to each of the following Celcius temperatures: -10, 0, 15, 100? McGraw Hill 6th edition, chaper 6, sec. 1, problem # 46. Thanks for your help!!!!
Answer by funmath(2933)

I accidentally lost this problem yesterday when I tried to answer it. I hope you're the same student. I'm sorry.
F=9/5(-10)+32
F=-90/5+32
F=-18+32
F=14 degrees F
--------------------
F=9/5(0)+32
F=0+32
F=32 degrees F
---------------------
F=9/5(15)+32
F=135/5+32
F=27+32
F=59 degrees F
----------------------
F=9/5(100)+32
F=900/5+32
F=180+32
F=212 degrees F
Happy Calculating!